### v2

# Visuomotor Policy Learning via Coarse-to-Fine Autoregressive Prediction

Abstract
In robotic visuomotor policy learning, diffusion-based models have achieved significant success in improving the accuracy of action trajectory generation compared to traditional autoregressive models. However, they suffer from inefficiency due to multiple denoising steps and limited flexibility from complex constraints. In this paper, we introduce Coarse-to-Fine AutoRegressive Policy (CARP), a novel paradigm for visuomotor policy learning that redefines the autoregressive action generation process as a coarse-to-fine, next-scale approach. CARP decouples action generation into two stages: first, an action autoencoder learns multi-scale representations of the entire action sequence; then, a GPT-style transformer refines the sequence prediction through a coarse-to-fine autoregressive process. This straightforward and intuitive approach produces highly accurate and smooth actions, matching or even surpassing the performance of diffusion-based policies while maintaining efficiency on par with autoregressive policies. We conduct extensive evaluations across diverse settings, including single-task and multi-task scenarios on state-based and image-based simulation benchmarks, as well as real-world tasks. CARP achieves competitive success rates, with up to a 10% improvement, and delivers 10× faster inference compared to state-of-the-art policies, establishing a high-performance, efficient, and flexible paradigm for action generation in robotic tasks.
1 Introduction
Figure 1: Policy Comparison. The representative performance among Behavior Transformer [52] served as an autoregressive policy, Diffusion Policy [11], and our approach in the state-based Robomimic square task experiment. CARP achieves a superior balance of performance and efficiency.
(a) Autoregressive Policy
(b) Diffusion Policy
(c) CARP (Ours)
Figure 2: Structure of Current Policies. $\hat{\boldsymbol{a}}$ is the predicted action, $\boldsymbol{a}^{k}$ denotes the refining action at step $k$, $\boldsymbol{s}$ is the historical condition. a) Autoregressive Policy predicts the action step-by-step in the next-token paradigm. b) Diffusion Policy models the noise process used to refine the action sequence. c) CARP refines action sequence predictions autoregressively from coarse to fine granularity.
Policy learning from demonstrations, formulated as the supervised regression task of mapping observations to actions, has proven highly effective across various robotic tasks, even in its simplest form. Replacing the policies with generative models, particularly those stemming from the vision community, has opened new avenues for improving performance, enabling robots to achieve the high precision required for complex tasks.
Existing approaches have explored different generative modeling techniques to address challenges in visuomotor policy learning. Autoregressive Modeling (AM) [52, 70, 12, 30] provides a straightforward and efficient out-of-the-box solution, benefitting from its scalability, flexibility, and mature exploration at lower computation requirements. However, AM’s next-token prediction paradigm often fails to capture long-range dependencies, global structure, and temporal coherence [26], leading to poor performance, which is essential for many robotic tasks. Recently, Diffusion Modeling (DM) [11, 61, 51] has emerged as a promising alternative, bridging the precision gap in AM by modeling the gradient of the action score function to learn multimodal distributions. Nevertheless, DM requires multiple steps of sequential denoising, making it computationally prohibitive for robotic tasks, especially for robotic tasks requiring efficient real-time inference in on-board compute-constrained environments. Additionally, DDPM [19]’s rigid generative process lacks flexibility and adaptability for tasks with long-term dependencies, often leading to cumulative errors and reduced robustness over extended time spans.
Both AM and DM have their respective advantages and limitations, which are often orthogonal and difficult to balance in practical applications. In this work, we aim to resolve this trade-off by introducing a novel generative paradigm for robot visuomotor policy learning that predicts entire action sequences from a coarse-to-fine granularity in a next-scale prediction framework. This approach allows our model to achieve performance levels comparable to DM while achieving AM’s efficiency and flexibility.
Our primary contribution is the introduction of a Coarse-to-Fine AutoRegressive Policy (CARP), a hybrid framework that combines AM’s efficiency with DM’s high performance to meet the demands of real-world robotic manipulation. Specifically, our contributions are as follows:
  • Multi-scale action tokenization: We propose a multi-scale tokenization method for action sequences that captures the global structure and maintains temporal locality, effectively addressing AM’s myopic limitations.
  • Coarse-to-fine autoregressive prediction: This mechanism refines action sequences in the latent space using Cross-Entropy loss with relaxed Markovian assumptions during iterations, achieving DM-like performance with high efficiency.
  • Comprehensive sim & real experiments: Extensive experiments demonstrate CARP’s effectiveness in both simulated and real-world robotic manipulation tasks.
In summary, we present CARP, a novel visuomotor policy framework that synergizes the strengths of AM and DM, offering high performance, efficiency, and flexibility.
(a) Multi-Scale Action Tokenization
(b) Coarse-to-Fine Autoregressive Prediction
Figure 3: Overview of the Two Stages of CARP. a) A multi-scale action autoencoder extracts token maps $\boldsymbol{r}_{1},\boldsymbol{r}_{2},\dots,\boldsymbol{r}_{K}$ to represent the action sequence at different scales, trained using the standard VQVAE loss. b) The autoregressive prediction is reformulated as a coarse-to-fine, next-scale paradigm. The sequence is progressively refined from coarse token map $\boldsymbol{r}_{1}$ to finer granularity token map $\boldsymbol{r}_{K}$, where each $\boldsymbol{r}_{k}$ contains $l_{k}$ tokens. An attention mask ensures that each $\boldsymbol{r}_{k}$ attends only to the preceding $\boldsymbol{r}_{1:k-1}$ during training. A standard Cross-Entropy loss is used for training. During inference, the token maps $\boldsymbol{r}_{1:K}$ are collectively decoded into continuous actions for execution.
2 Background
We start by introducing the background, focusing on three key areas: problem formulation, conventional autoregressive policies, and diffusion-based policies.
2.1 Problem Formulation
Problem formulation will consider a task $\mathcal{T}$, where there are $N$ expert demonstrations $\{\tau_{i}\}_{i=1}^{N}$. Each demonstration $\tau_{i}$ is a sequence of state-action pairs. We formulate robot imitation learning as an action sequence prediction problem [46, 11, 61], training a model to minimize the error in future actions conditioned on historical states. Specifically, imitation learning minimize the behavior cloning loss $\mathcal{L}_{bc}$ formulated as
$\mathcal{L}_{bc}=\mathbb{E}_{\boldsymbol{s},\boldsymbol{a}\sim\mathcal{T}}\left[\sum_{t=0}^{T}\mathcal{L}\left(\pi_{\boldsymbol{\theta}}(\boldsymbol{a}_{H}|\boldsymbol{s}_{O}),\boldsymbol{a}_{H}\right)\right],$ (1)
where $\boldsymbol{a}$ represents the action, $\boldsymbol{s}$ denotes the state or observation according to the specific task description, $t$ is the current time step, $H$ is the prediction horizon, and $O$ is the historical horizon. For notational simplicity, we denote the action sequence $\boldsymbol{a}_{t:t+H-1}$ as $\boldsymbol{a}_{H}$ and the state sequence $\boldsymbol{s}_{t-O+1:t}$ as $\boldsymbol{s}_{O}$. Here, $\mathcal{L}_{bc}$ represents a supervised action prediction loss, such as mean squared error or negative log-likelihood, $T$ is the length of the demonstration, and $\theta$ represents the learnable parameters of the policy network $\pi_{\boldsymbol{\theta}}$.
2.2 Autoregressive Policy
Autoregressive policies leverage the efficiency and flexibility of autoregressive models (GPT-style Decoders). Recent advancements, such as action chunking [52, 70], redefine this paradigm as a multi-token, one-pass prediction method (see Appendix B for further details). We refer to this approach as Autoregressive Policy (AP). In this approach, the next-token posits the probability of observing the current action $\boldsymbol{a}_{t}$ depends solely on its previous states $\boldsymbol{s}_{O}$, which allows for the factorization of the likelihood of sequence with length $H$ as:
$p\left(\boldsymbol{a}_{t},\boldsymbol{a}_{t+1},...,\boldsymbol{a}_{t+H-1}\right)=\prod^{t+H-1}_{k=t}p\left(\boldsymbol{a}_{k}|\boldsymbol{s}_{O}\right).$ (2)
However, it introduces several issues that hinder its performance. The linear, step-by-step unidirectional dependency of action prediction may overlook the global structure [26], making it challenging to capture long-range dependencies and holistic coherence or temporal locality [22] in complex scenes or sequences, which restricts their generalizability in tasks requiring bidirectional reasoning, as shown in Fig. 2(a). For example, it cannot predict the former actions given the near-terminal state.
2.3 Diffusion Policy
Diffusion-based policies [11] utilize Denoising Diffusion Probabilistic Models [19] to approximate the conditional distribution $p(\boldsymbol{a}_{H}|\boldsymbol{s}_{O})$, through modeling the noise during the denoising process from Gaussian noise to noise-free output as
$\boldsymbol{a}_{H}^{k+1}=\alpha(\boldsymbol{a}_{H}^{k}-\gamma\epsilon_{\theta}(\boldsymbol{s}_{O},\boldsymbol{a}_{H}^{k},k)+\mathcal{N}),$ (3)
where $\epsilon_{\theta}$ is the learnable noise network, $k$ is the current denosing step, $\mathcal{N}$ is the Gaussian noise, $\alpha$ and $\gamma$ are the hyper-parameters.
Diffusion-based policies show impressive performance in robotic manipulation tasks due to their action generation which gradually refines from random samples which we can abstract as a coarse-to-fine process. This kind of process alike human movement or natural thinking flows in line with human intuition. However, they suffer from poor convergence due to their design. To address this, multiple denoising steps are required, constrained by the Markovian assumption, where $\boldsymbol{a}_{H}^{k+1}$ depends only on the previous step $\boldsymbol{a}_{H}^{k}$ (as shown in Fig. 2(b)), which leads to significant runtime inefficiency. Moreover, they lack the ability to leverage generation context effectively, hindering scalability to complex scenes [16].
3 Method
To address the limitations of existing methods, we propose Coarse-to-Fine AutoRegressive Policy (CARP), a novel visuomotor policy framework that combines the high performance of recent diffusion-based policies with the sample efficiency and flexibility of traditional autoregressive policies.
CARP achieves these advantages through a redesigned autoregressive modeling strategy. Specifically, we shift from conventional next-token prediction to a coarse-to-fine, next-scale prediction approach, using multi-scale action representations. CARP’s training is divided into two stages: multi-scale action tokenization and coarse-to-fine autoregressive prediction, as shown in Fig. 3.
In this section, we first explain the construction of multi-scale action token maps, followed by the coarse-to-fine autoregressive prediction approach for action generation. Key implementation details are provided at the end.
3.1 Multi-Scale Action Tokenization
Instead of focusing on individual action steps, we extract representations at multiple scales across the entire action sequence. We propose a novel multi-scale action quantization autoencoder that encodes a sequence of actions into $K$ discrete token maps, $\boldsymbol{R}=(\boldsymbol{r}_{1},\boldsymbol{r}_{2},\dots,\boldsymbol{r}_{K})$, which are used for both training and inference. Our approach builds on the VQVAE architecture [60], incorporating a modified multi-scale quantization layer [58] to enable hierarchical encoding.
Encoder and Decoder. As illustrated in Fig. 3(a), the actions are first organized into an action sequence $\boldsymbol{A}^{H\times D}$, where $H$ denotes the prediction horizon and $D$ represents the dimensionality of the action $\boldsymbol{a}$. Given that each dimension in the action space is orthogonal to the others, and that there exists a natural temporal dependency within an action sequence, we employ a 1D temporal convolutional network (1D-CNN) along the time dimension, as used in [22], for both the encoder $\mathcal{E}(\cdot)$ and decoder $\mathcal{D}(\cdot)$. Let $\boldsymbol{F}=\mathcal{E}(\boldsymbol{A})\in\mathbb{R}^{L\times C}$, where $L$ denotes the compressed length of the temporal dimension with $L\leq H$, and $C$ represents the dimensionality of the feature map $\boldsymbol{F}$.
Quantization. We introduce a quantizer with a learnable codebook $\mathcal{Z}\in\mathbb{R}^{V\times C}$, containing $V$ code vectors, each of dimension $C$. The quantization process (lines 4-11 in Algorithm 1) generates the action sequence representation iteratively across multiple scales. At scale $k$, it produces action token map $\boldsymbol{r}_{k}$ to represent the sequence, which consists of $l_{k}$ tokens $q$. In our implementation, we set $l_{k}=k$. The function $\operatorname{Lookup}(\mathcal{Z},v)$ retrieves the $v$-th code vector from $\mathcal{Z}$. The quantization function is defined by $\boldsymbol{r}=\mathcal{Q}(\boldsymbol{F})$ as following:
$q=\operatorname*{arg\,min}_{v\in[V]}\operatorname{dist}\left(\operatorname{Lookup}(\mathcal{Z},v),\boldsymbol{f}\right)\in[V],$ (4)
where token $q$ represents the nearest vector in $\mathcal{Z}$ for a given feature vector $\boldsymbol{f}\in\mathbb{R}^{1\times C}$ in the feature map $\boldsymbol{F}$, based on a distance function $\operatorname{dist}(\cdot)$.
Specifically, we adopt a residual-style design [29, 58] for feature maps $\boldsymbol{F}$ and $\hat{\boldsymbol{F}}$, as shown in lines 4-11 of Algorithm 1. This design ensures that each finer-scale representation $\boldsymbol{r}_{k}$ depends only on its coarser-scale predecessors $(\boldsymbol{r}_{1},\boldsymbol{r}_{2},\dots,\boldsymbol{r}_{k-1})$, facilitating a multi-scale representation. A shared codebook $\mathcal{Z}$ is used across all scales, ensuring that multi-scale token maps $\boldsymbol{R}=(\boldsymbol{r}_{1},\boldsymbol{r}_{2},\dots,\boldsymbol{r}_{K})$, where $K$ is the number of scales, are drawn from a consistent vocabulary $[V]$. To preserve information during upsampling, we employ $K$ additional 1D convolutional layers $\{\phi_{k}\}_{k=1}^{K}$ [58], as illustrated in lines 9-10 of Algorithm 1.
Loss. The final approximation $\hat{\boldsymbol{F}}$ of the original feature map $\boldsymbol{F}$ is combined residually with the multi-scale representation derived from the codebook $\mathcal{Z}$, based on each $\boldsymbol{r}$. The reconstructed action sequence is then obtained as $\hat{\boldsymbol{A}}=\mathcal{D}(\hat{\boldsymbol{F}})$. To train the quantized autoencoder, a typical VQVAE loss $\mathcal{L}$ is minimized as following
$\mathcal{L}=\underbrace{\|\boldsymbol{A}-\hat{\boldsymbol{A}}\|_{2}}_{\mathcal{L}_{\text{recon}}}+\underbrace{\|\text{sg}(\boldsymbol{F})-\hat{\boldsymbol{F}}\|_{2}}_{\mathcal{L}_{\text{quant}}}+\underbrace{\|\boldsymbol{F}-\text{sg}(\hat{\boldsymbol{F}})\|_{2}}_{\mathcal{L}_{\text{commit}}},$ (5)
where $\mathcal{L}_{\text{recon}}$ minimizes the difference between the original action sequence $\boldsymbol{A}$ and the reconstruction $\hat{\boldsymbol{A}}$. $\text{sg}(\cdot)$ denotes stop-gradient. $\mathcal{L}_{\text{quant}}$ aligns the quantized feature map $\hat{\boldsymbol{F}}$ with the original $\boldsymbol{F}$, and $\mathcal{L}_{\text{commit}}$ encourages the encoder to commit to codebook entries, preventing codebook collapse. For $\mathcal{L}_{\text{quant}}$ and $\mathcal{L}_{\text{commit}}$, we calculate every residual calculating moment of each scale $\boldsymbol{r}_{k}$ as in Algorithm 1. After training, the autoencoder $\{\mathcal{E},\mathcal{Q},\mathcal{D}\}$ tokenizes actions for subsequent coarse-to-fine autoregressive modeling.
Discussion. The tokenization strategy described above allows the multi-scale tokens $\boldsymbol{R}$, extracted from the action sequence $\boldsymbol{A}\in\mathbb{R}^{H\times D}$ via temporal 1D convolutions, to inherently preserve temporal locality [22]. Additionally, the hierarchical extraction captures the global structure, enabling the model to treat the action sequence as a unified entity.
Unlike traditional autoregressive policies that predict each action token independently (see Eq. 2 and Fig. 2(a)), CARP leverages dual capabilities to capture both local temporal dependencies and global structure across the entire action sequence. This approach produces smoother transitions and more stable action sequences. Through multi-scale encoding, CARP overcomes the short-sighted limitations of conventional autoregressive models, yielding more robust, coherent, and precise behaviors over extended time horizons.
Algorithm 1 Multi-Scale Action VQVAE
1:  Inputs: Action sequence $\boldsymbol{A}$
2:  Hyperparameters: Number of scales $K$, length of each scale $(l_{k})_{k=1}^{K}$ , length of feature map’s temporal dimension $L$
3:  Initialize: $\boldsymbol{F}\leftarrow\mathcal{E}(\boldsymbol{A})$, $\hat{\boldsymbol{F}}\leftarrow 0$, $R\leftarrow[]$
4:  for $k=1$ to $K$ do
5:     $\boldsymbol{r}_{k}\leftarrow\mathcal{Q}(\text{Interpolate}(\boldsymbol{F},l_{k}))$
6:     $\boldsymbol{R}\leftarrow\boldsymbol{R}\cup\{\boldsymbol{r}_{k}\}$
7:     $\boldsymbol{Z}_{k}\leftarrow\text{Lookup}(\mathcal{Z},\boldsymbol{r}_{k})$
8:     $\boldsymbol{Z}_{k}\leftarrow\text{Interpolate}(\boldsymbol{Z}_{k},L)$
9:     $\boldsymbol{F}\leftarrow\boldsymbol{F}-\phi_{k}(\boldsymbol{Z}_{k})$
10:     $\hat{\boldsymbol{F}}\leftarrow\hat{\boldsymbol{F}}+\phi_{k}(\boldsymbol{Z}_{k})$
11:  end for
12:  $\hat{\boldsymbol{A}}\leftarrow\mathcal{D}(\hat{\boldsymbol{F}})$
13:  Return: Multi-scale token maps $\boldsymbol{R}$, reconstructed action sequence $\hat{\boldsymbol{A}}$
3.2 Coarse-to-Fine Autoregressive Prediction
Using multi-scale action sequence representations, we shift from traditional next-token prediction to a next-scale prediction approach, progressing from coarse to fine granularity.
Prediction. As we receive the multi-scale representation tokens $(\boldsymbol{r}_{1},\boldsymbol{r}_{2},\dots,\boldsymbol{r}_{K})$, each $\boldsymbol{r}$ represents the same action sequence in different scales. The autoregressive likelihood can be formulated as
$p(\boldsymbol{r}_{1},\boldsymbol{r}_{2},\dots,\boldsymbol{r}_{K})=\prod_{k=1}^{K}p(\boldsymbol{r}_{k}\mid\boldsymbol{r}_{1},\boldsymbol{r}_{2},\dots,\boldsymbol{r}_{k-1};\boldsymbol{s}_{O}),$ (6)
where each autoregressive unit $\boldsymbol{r}_{k}\in[V]^{l_{k}}$ is the token map at scale $k$ containing $l_{k}$ tokens $q$. The predix sequence $(\boldsymbol{r}_{1},\boldsymbol{r}_{2},\dots,\boldsymbol{r}_{k-1})$ is served as the condition for $\boldsymbol{r}_{k}$, accompanying with the historical state sequence $\boldsymbol{s}_{O}$. This kind of next-scale prediction methodology is what we define as coarse-to-fine autoregressive prediction.
Due to the residual-style quantization, during autoregressive prediction, we first embed the previous scale token map $\boldsymbol{r}_{k-1}$ and then reconstruct the next scale feature map $\boldsymbol{e}_{k}$, as shown in Fig. 3(b). This feature map $\boldsymbol{e}_{k}$ is then used as input for predicting the next scale token map $\boldsymbol{r}_{k}$. The token maps $\boldsymbol{r}_{1:K}$ are decoded by the multi-scale autoencoder into continuous actions, following the residual-style process.
Loss. During the $k$-th autoregressive step, all distributions over the $l_{k}$ tokens in $\boldsymbol{r}_{k}$ will be generated in parallel, with the coarse-to-fine dependency ensured by a block-wise causal attention mask [58]. To optimize the autoregressive model, we utilize the standard Cross-Entropy loss to capture the difference between the predicted token map $\hat{\boldsymbol{r}}$ and the token map $\boldsymbol{r}$ from the ground truth action sequence:
$\mathcal{L}_{\text{Cross-Entropy}}=\sum_{k=1}^{K}\sum_{i=1}^{l_{k}}\left[-\sum_{v=1}^{V}\boldsymbol{r}_{k}^{i,v}\log\hat{\boldsymbol{r}}_{k}^{i,v}\right],$ (7)
where $l_{k}$ is the length of each scale, $V$ is the action vocabulary size, and $K$ is the number of scales.
Discussion. CARP models the entire trajectory holistically, progressively refining actions from high-level intentions to fine-grained details (see Fig. 2(c)). This approach aligns more closely with natural human behavior, where movement is guided by overarching intentions rather than step-by-step planning.
Instead of the commonly used MSE loss [70, 59, 5], CARP employs Cross-Entropy loss (Eq. 7) in behavior cloning, as MSE tends to enforce a unimodal distribution that can be detrimental to manipulation tasks [52]. CARP’s iterative refinement process resembles the denoising steps in diffusion models to achieve high accuracy.
Furthermore, our approach models actions directly rather than modeling noise, enabling faster convergence in low-dimensional manifolds [36, 64]. By operating in the latent space with action sequence tokens, CARP mitigates trajectory anomalies that can disrupt prediction, unlike methods that work on raw actions [11]. This latent-space representation allows CARP to focus on the essential components of actions, resulting in smoother and more efficient predictions. In contrast, traditional diffusion models rely on Markovian processes, which compress all information into progressively noisier inputs from previous levels, often hindering efficient learning and requiring more inference steps [16] (see Eq. 3 and Fig. 2(b)). CARP relaxes this constraint by allowing each scale $\boldsymbol{r}_{k}$ to depend on all prior scales $\boldsymbol{r}_{1:k-1}$ instead of the only previous scale $\boldsymbol{r}_{k-1}$. This structure enables CARP to generate high-quality trajectories with significantly fewer steps.
3.3 Implementation Details
Our primary focus is on the CARP algorithm, and we adopt a simple model architecture design.
Tokenization. Due to the disentangling of the action prediction, imprecise coarse-to-fine actions representation can decrease the upper limit of model performance. Considering the discontinuity of the most representation for rotation space like Euler angle or quaternion will increase the unstable training process, we utilize rotation6d [71] to get stable multi-scale tokens. And for the distance function $\operatorname{dist}(\cdot)$, we use cosine similarity rather than Euclidean distance which is the cause of unstable training based on our observations. In practice, due to the orthogonality between dimensions, we use a separate VQVAE [58] for each dimension of the action space, to gain stable training. All convolutions we used in CARP are 1D to capture the time-dimension features.
Autoregressive. We adopt the architecture of standard decoder-only transformers akin to GPT-2 [58, 9]. The state $\boldsymbol{s}_{O}$ is used to generate the initial coarse-scale token map and then is utilized as adaptive normalization [40] for the subsequent predictions. During training, we observe that incorporating Exponential Moving Average (EMA) [18] enhances both training stability and performance, yielding a 4-5% improvement, consistent with findings in [11]. During the inference, kv-caching can be used and no mask is needed, as shown in Fig. 10.
4 Experiment
In this section, we evaluate CARP on diverse robotics tasks, including state-based and image-based benchmarks in single-task and multi-task settings. CARP’s performance is assessed based on task success rates, inference speed, and model parameter scale. Additionally, we validate CARP’s practical effectiveness by deploying it on real-world tasks using a UR5e robotic arm, comparing its performance against state-of-the-art diffusion-based policies.
Our experiments are structured to address the following key research questions:
Figure 4: Single-Task Simulation Setup. We evaluate three tasks from the Robomimic [38] benchmark—Lift, Can, and Square—ordered by increasing difficulty, along with a Kitchen task [17] on the far left.
  • RQ1: Can CARP match the accuracy and robustness of current state-of-the-art diffusion-based policies?
  • RQ2: Does CARP maintain high inference efficiency and achieve fast convergence?
  • RQ3: Does CARP leverage the flexibility benefits of a GPT-style architecture?
4.1 Evaluation on Simulation Benchmark
We first evaluate CARP on a set of simulated tasks commonly used to benchmark diffusion-based models, assessing its ability to match their performance while significantly improving computational efficiency.
Policy Lift Can Square Params/M Speed/s
BET [52] 0.96 0.88 0.54 0.27 2.12
DP-C [11] 1.00 0.94 0.94 65.88 35.21
DP-T [11] 1.00 1.00 0.88 8.97 37.83
CARP (Ours) 1.00 1.00 0.98 0.65 3.07
Table 1: State-Based Simulation Results (State Policy). We report the average success rate of the top 3 checkpoints, along with model parameter scales and inference time for generating 400 actions. CARP significantly outperforms BET and achieves competitive performance with state-of-the-art diffusion models, while also surpassing DP in terms of model size and inference speed.
Policy Lift Can Square Params/M Speed/s
IBC [15] 0.72 0.02 0.00 3.44 32.35
DP-C [11] 1.00 0.97 0.92 255.61 47.37
DP-T [11] 1.00 0.98 0.86 9.01 45.12
CARP (Ours) 1.00 0.98 0.88 7.58 4.83
Table 2: Image-Based Simulation Results (Visual Policy). Results show that CARP consistently balances high performance and high efficiency. We highlight our results in light-blue.
Experimental Setup. We use the Robomimic [38] benchmark suite, which is widely adopted for evaluating diffusion-based models [11, 61, 42]. We evaluate CARP on both state-based and image-based datasets collected from expert demonstrations, with each task containing 200 demonstrations. The selected tasks are standard in single-task evaluations, as shown in Fig. 4 and Fig. 12. For evaluating the ability to learn multiple long-horizon tasks, we utilize the Franka Kitchen environment [17] which contains 7 objects for interaction and comes with a human demonstration dataset of 566 demonstrations, each completing 4 tasks in arbitrary order (see  Fig. 14 for details).
Baselines. We compare CARP with previous autoregressive policies as well as recent diffusion policies. Behavior Transformer (BET) [52] is an autoregressive policy with action discretization and correction mechanisms, similar to offset-based prediction. Implicit Behavior Cloning (IBC) [15] utilizes energy-based models for supervised robotic behavior learning. Diffusion Policy (DP) [11] combines a denoising process with action prediction, implemented in two variants: CNN-based (DP-C) and Transformer-based (DP-T).
Policy p1 p2 p3 p4 Params Speed
BET [52] 0.96 0.84 0.6 0.20 0.30 1.95
DP-C [11] 1.00 1.00 1.00 0.96 66.94 56.14
DP-T [11] 1.00 0.99 0.98 0.96 80.42 56.32
CARP (Ours) 1.00 1.00 0.98 0.98 3.88 2.01
[…]
Metrics. For each task, we evaluate the policies by running 50 trials with random initializations to compute the success rate and report the average success rate across the top 3 checkpoints. For the Kitchen task, success rates are reported with increasing difficulty, determined by the number of interactive objects (from p1 to p4). Inference speed is measured on an A100 GPU by averaging the time required for 400 action predictions (280 for the Kitchen task) over 5 runs to ensure robustness. We also record the parameter count for each model using the same PyTorch implementation interface.
Figure 5: Visualization of the Trajectory and Refining Process. The left panel shows the final predicted trajectories for each task, with CARP producing smoother and more consistent paths than Diffusion Policy (DP). The right panel visualizes intermediate trajectories during the refinement process for CARP (top-right) and DP (bottom-right). DP displays considerable redundancy, resulting in slower processing and unstable training, as illustrated by 6 selected steps among 100 denoising steps. In contrast, CARP achieves efficient trajectory refinement across all 4 scales, with each step contributing meaningful updates.
Implementation Details. For baseline models, we follow the same implementation and training configurations provided by [11]. In state-based experiments, we set the observation horizon $O=2$ and the prediction horizon $H=16$ across all models. For image-based experiments, we set $O=1$ and $H=16$ for better transferability to real-world scenarios. As per the benchmark, only the first 8 actions in the prediction horizon are executed starting from the current step (see Appendix C). For CARP, we first train an action VQVAE model (see Sec. 3.1) following [29], using $V=512$, $C=8$, a batch size of 256, and 300 epochs per task Given a horizon $H=16$, we design multi-scale representations with scales of $1,2,3$ and $4$ to capture coarse-to-fine information across the action sequence. We then train an autoregressive GPT-2 style, decoder-only transformer (see Sec. 3.2), based on [58], using the same training settings as the benchmark, with a batch size of 256 for state-based experiments (4000 epochs) and a batch size of 64 for image-based experiments (3000 epochs).
Results. As demonstrated in Tables 1, 2, and 3, CARP consistently achieves comparable performance to state-of-the-art diffusion models across both state-based and image-based tasks, answering RQ1. Notably, CARP significantly outperforms diffusion policies in terms of inference speed, being approximately 10 times faster, with only 1-5% of the parameters required by diffusion models, thereby strongly supporting RQ2 regarding CARP’s efficiency.
Analysis. To further examine CARP’s stability and efficiency, we visualize spatial trajectories along the $xyz$ axes for the Can and Square tasks in Fig. 5. In each task’s left panel, CARP consistently reaches specific regions (light grey area) to perform task-related actions, such as positioning for object grasping or placement. Compared to diffusion-based models, CARP’s trajectories are smoother and more consistent, underscoring its stability and accuracy (supporting RQ1). The right panels of Fig. 5 compare CARP’s coarse-to-fine action predictions with 6 selected denoising steps of the diffusion model. CARP achieves accurate predictions within just 4 coarse-to-fine steps, whereas the diffusion model requires numerous denoising iterations, with many early steps introducing redundant computations. This analysis further supports RQ2 and highlights CARP’s efficiency and fast convergence advantage over diffusion policies in generating accurate actions with fewer refinement steps, as detailed in Sec. 3.2.
4.2 Evaluation on Multi-Task Benchmark
We evaluate CARP on the MimicGen [39] multi-task simulation benchmark, widely used by the state-of-the-art Sparse Diffusion Policy (SDP) [61], to demonstrate CARP’s flexibility, a result of its GPT-style autoregressive design.
Figure 6: Multi-Task Simulation Setup. We evaluate eight tasks from the MimicGen [39] benchmark: Coffee, Hammer Cleanup, Mug Cleanup, Nut Assembly, Square, Stack, Stack Three, and Threading, listed left-to-right and top-to-bottom.
Policy Prams/M Speed/s Coffee Hammer Mug Nut Square Stack Stack three Threading Avg.
TCD [34] 156.11 107.15 0.77 0.92 0.53 0.44 0.63 0.95 0.62 0.56 0.68
SDP [61] 159.85 112.39 0.82 1.00 0.62 0.54 0.82 0.96 0.80 0.70 0.78
CARP (Ours) 16.08 6.92 0.86 0.98 0.74 0.78 0.90 1.00 0.82 0.70 0.85
Table 4: Multi-Task Simulation Results (Visual Policy). Success rates are averaged across the top three checkpoints for each task, as well as the overall average across all tasks. We also report parameter count and inference time for generating 400 actions. CARP outperforms diffusion-based policies by 9%-25% in average performance, with significantly fewer parameters and over 10× faster inference.
Experimental Setup. MimicGen extends benchmark Robomimic [38] by including 1K–10K human demonstrations per task, with diverse initial state distributions for enhanced generalization in multi-task evaluation. It consists of 12 robosuite [72] tasks powered by MuJoCo and 4 high-precision tasks from Isaac Gym Factory. We select 8 robosuite tasks—Coffee, Hammer, Mug, Nut, Square, Stack, Stack Three, and Threading—for evaluation, each with 1K training trajectories, following the settings in SDP [61]. Visualization is shown in Fig. 6 and Fig. 13.
Baselines. We compare CARP with two baselines: Task-Conditioned Diffusion (TCD) [2, 34]: A basic diffusion-based multi-task policy and Sparse Diffusion Policy (SDP) [61]: A transformer-based diffusion policy that leverages Mixture of Experts (MoE) [53]. Both baselines are trained using visual inputs.
Metrics. Success rates are reported for each task as the average of the best three checkpoints. For Nut Assembly, partial success (e.g., placing one block inside the cylinder) is assigned a score of 0.5, while full success is scored as 1. For all other tasks, a score of 1 is awarded only when strict success criteria are fully satisfied. Additionally, we calculate the average success rate across all tasks. To evaluate model efficiency, we report both the parameter scale calculated by the Pytorch interface and the inference time required to predict 400 actions on a single A100 GPU.
Implementation Details. For CARP, we extend the single-task implementation by incorporating a task embedding as an additional condition, alongside the observation sequence $\boldsymbol{s}$. We also use a moderately deeper decoder-only transformer in GPT-2 style. CARP is trained with a batch size of 512 for 200 epochs on an A100 GPU. Baseline models follow the same training settings as SDP [61]. This minimal modification enables CARP to adapt to multi-task learning seamlessly.
Results. As shown in Tab. 4, CARP achieves up to a 25% average improvement in success rates compared to state-of-the-art diffusion-based policies, highlighting its strong performance. With minimal modification, CARP seamlessly transitions from single-task to multi-task learning, further demonstrating its flexibility, a benefit of its GPT-style architecture. Additionally, as shown in Tab. 4, CARP achieves over 10× faster inference speed and uses only 10% of the parameters compared to SDP. Leveraging its GPT-style autoregressive design [7, 47], CARP can seamlessly adapt to multi-task settings without complex design modifications [61], demonstrating the flexibility of this architecture, which strongly supports RQ3. These results strongly demonstrate CARP’s flexibility, solidifying its role as a high-performance and high-efficiency approach for visuomotor robotic policies.
4.3 Evaluation on Real-World
In this section, we evaluate our approach, CARP, on real-world tasks under compute-constrained conditions, comparing its performance and efficiency against baseline methods.
Figure 7: Real-World Setup. The left panel shows the environment used for the experiment and demonstration collection. The right panel shows the trajectory from the Cup and Bowl datasets.
Experimental Setup. To validate CARP’s real-world applicability, we design two manipulation tasks:
  1. 1.
    Cup: The robot must locate a cup on the table, pick it up, move to the right area of the table, and put it down steadily (top-right, Fig. 7).
  2. 2.
    Bowl: The robot needs to identify a smaller bowl and a larger pot on the table, pick up the bowl, and place it inside the pot (bottom-right, Fig. 7).
We use a UR5e robotic arm with a Robotiq-2f-85 gripper, equipped with two RGB cameras: one mounted on the wrist and one in a third-person perspective (left panel, Fig. 7). The robot is controlled through 6D end-effector positioning, with inverse kinematics for joint angle calculation. For teleoperation, we collected 50 human demonstration trajectories for each task using a 3D Connexion space mouse.
Baselines. As a baseline, we reproduce the CNN-based image-based diffusion policy [11], adapting the model’s input size to accommodate our observational setup.
Figure 8: Real-World Results (Visual Policy). We report the average success rate across 20 trials and the inference speed as action prediction frequency. CARP achieves competitive success rates with significantly faster inference compared to DP.
Metrics. For each trained policy, we report the average success rate across 20 trials per task, with the initial positions randomized. We also measure inference speed on an NVIDIA GeForce RTX 2060 GPU, reporting action prediction frequency in Hertz.
Implement Details. For both models, the input consists of current visual observations from the wrist and scene cameras (resolution: $120\times 160$), as well as proprioceptive data from the robotic arm. We execute 8 predicted actions out of a horizon of 16 predictions. We train the diffusion policy for 3000 epochs with a batch size of 64. For CARP, we use the same visual policy structure as in the simulation tasks, training the action VQVAE for 300 epochs with a batch size of 256, and the decoder-only transformer for 3000 epochs with a batch size of 64.
Results. As shown in Fig. 8 and Fig. 9, CARP achieves comparable or superior performance, with up to a 10% improvement in success rate over the diffusion policy across all real-world tasks, supporting RQ1. Additionally, CARP achieves approximately 8×8$\times$8 × faster inference than the baseline on limited computational resources, demonstrating its suitability for real-time robotic applications, thus supporting RQ2 (see Fig. 15 for further visualization).
Figure 9: Visualization of CAPR on Real-World Tasks. CARP generates smooth and successful trajectories for the Cup and Bowl tasks, with temporal progression from left to right.
5 Related Work
5.1 Visual Generation
Advancements in generative models for visual generation have significantly influenced the robotics community. Autoregressive models generate images in a raster-scan fashion using discrete tokens from image tokenizers [14, 20, 14, 60]. GPT-2-style transformers [29, 4, 48, 63] demonstrate strong performance by generating tokens sequentially. Recent work has scaled these models to billions of parameters, achieving impressive text-to-image synthesis results [69, 35] and robotic action generation [13, 54]. VAR [58] introduces a new next-scale autoregressive paradigm that shifts image representation from patches to scales. This framework [58], has been applied across tasks [33, 32, 67, 43, 66, 37, 16]. Studies [58, 56] show that autoregressive models can surpass diffusion models in achieving state-of-the-art performance, serving as a key inspiration for our work.
5.2 Visuomotor Policy Learning
Behavior cloning [8] has proven effective, especially in autonomous driving [41] and manipulation [68], offering a simpler alternative to complex reinforcement learning. Explicit policies directly map states or observations to actions [68], enabling efficient inference. However, they struggle with complex tasks. Techniques like action space discretization [65] and Mixture Density Networks (MDNs) [52, 12] have been proposed, though they face issues with exponential action space growth and hyperparameter sensitivity. Implicit policies, often based on Energy-Based Models (EBMs) [23, 15], offer flexibility but are challenging to train due to optimization instabilities. Diffusion models have been applied to robotic policy learning [11, 50, 2], showing their effectiveness for decision-making tasks. However, their multi-step denoising process is computationally expensive. Recent efforts focus on improving generalization [62], adapting to 3D environments [61, 64], and enhancing modularity via Mixture of Experts (MoE)  [61]. Consistency models [55, 36] accelerate inference but often compromise action prediction accuracy, along with inflexible model design.
6 Conclusion
In this work, we introduce Coarse-to-Fine Autoregressive Policy (CARP), a novel paradigm for robotic visuomotor policy learning that combines the efficiency of autoregressive modeling (AM) with the high performance of diffusion modeling (DM). CARP incorporates: 1) multi-scale action tokenization to capture global structure and temporal locality, addressing AM’s limitations in long-term dependency; 2) coarse-to-fine autoregressive prediction that refines actions from high-level intentions to detailed execution, achieving DM-like performance with AM-level efficiency through latent space prediction and relaxed Markovian constraints. The comprehensive evaluations, from simulation to real-world, demonstrate CARP’s effectiveness in balancing high performance, efficiency, and flexibility.
We hope this work will inspire further exploration into next-generation policy learning by leveraging GPT-style autoregressive models, advancing a more unified perspective on current generative modeling techniques (see Appx.A).
Supplementary Material
Appendix A Limitations and Future Work
In this work, we propose CARP, a next-generation paradigm for robotic visuomotor policy learning, which effectively balances the long-standing trade-off between high performance and high efficiency seen in previous autoregressive modeling (AM) and diffusion modeling (DM) approaches. Despite these advancements, there remain several limitations and opportunities for improvement in future research.
First, the model structure of CARP can be further optimized for simplicity. Currently, CARP employs a two-stage design, where the first stage utilizes separate multi-scale action VQVAE modules for each action dimension to address their orthogonality. Future work could focus on developing a unified one-stage method that integrates multi-scale tokenization with the coarse-to-fine prediction process, resulting in a more efficient and streamlined framework without compromising performance.
Second, the multimodal capabilities of CARP remain underdeveloped. To mitigate the unimodality issue inherent in traditional autoregressive policies trained with MSE loss, CARP adopts a cross-entropy loss, which retains the potential for multimodality. Compared to the Diffusion Policy’s multimodal ability [11] likely coming from DDPM’s integration over a Stochastic Differential Equation [19], CARP’s approach is relatively straightforward. Further enhancements are needed to fully realize and leverage its multimodal potential.
Third, CARP’s adoption of the GPT-style paradigm opens up promising yet unexplored possibilities. Beyond the flexibility already demonstrated, the contextual understanding capabilities inherent in GPT-style architectures [44] suggest that CARP could be extended to support multimodal inputs [3, 45] like tactile and auditory information and address robotic tasks requiring long-term dependency reasoning [1]. Additionally, its potential for generalization in few-shot and zero-shot learning scenarios [7] represents a particularly valuable direction for future exploration.
Finally, but not exhaustively, the scaling properties of CARP represent a promising frontier. The scaling laws established in existing GPT-style models [25] could be seamlessly applied to CARP, suggesting that increasing network capacity and leveraging larger pre-training datasets could lead to substantial performance gains. Furthermore, recent advances in Vision-Language-Action (VLA) [6, 27, 31] models present an opportunity to integrate CARP into such frameworks, which could further validate its scalability and capabilities.
Figure 10: Coarse-to-Fine Autoregressive Inference. During inference, we leverage kv-caching to enable coarse-to-fine prediction without the need for causal masks. The full set of token maps, $\boldsymbol{r}_{1:K}$, is collectively decoded by the action VQVAE into executable actions for the robotic arm.
Appendix B Definition of Autoregressive Policy
Figure 11: Conventional Autoregressive Policy. In reinforcement learning, conventional autoregressive policies generate action tokens sequentially, where each token is predicted based on the previously generated tokens. This differs from the action chunking prediction shown in Fig. 2(a).
Autoregressive policies naturally capitalize on the efficiency and flexibility of autoregressive models. Initial works from a reinforcement learning perspective applied models like Transformers to predict the next action using states or rewards as inputs [10, 21], as shown in Fig. 11 and formalized as:
$p\left(\boldsymbol{a}_{t},\boldsymbol{a}_{t+1},...,\boldsymbol{a}_{t+H-1}\right)=\prod^{t+H-1}_{k=t}p\left(\boldsymbol{a}_{k}|\boldsymbol{a}_{k-H:k-1},\boldsymbol{s}_{k-H:k}\right),$ (8)
where $\boldsymbol{s}_{k-H:k-1}$ represents the states or observations corresponding to the previous actions $\boldsymbol{a}_{k-H:k-1}$, and $\boldsymbol{s}_{k}$ is the current state or observation. Following this paradigm, several subsequent works [49, 24, 57] employ autoregressive models to predict one action at a time during inference.
More recently, the concept of action chunking [28], derived from neuroscience, has demonstrated notable benefits for imitation learning [52, 12, 70, 30]. In action chunking, individual actions are grouped and executed as cohesive units, leading to improved efficiency in storage and execution, as depicted in Fig. 2(a). This paradigm extends the capabilities of GPT-style decoders by modifying them to generate chunks of actions in one forward pass, replacing the traditional single-step autoregressive operation with a multi-token, pseudo-autoregressive process. The action generation process for this paradigm is described as:
$p\left(\boldsymbol{a}_{t},\boldsymbol{a}_{t+1},\dots,\boldsymbol{a}_{t+H-1}\right)=\prod^{t+H-1}_{k=t}p\left(\boldsymbol{a}_{k}|\boldsymbol{s}_{O}\right),$ (9)
where $O$ is the historical horizon. The model predicts the entire action sequence in one forward pass without strictly adhering to step-by-step autoregressive operations.
Given the significant performance improvements enabled by action chunking, we adopt this multi-token, one-forward-pass framework throughout the article and experiments when referring to autoregressive policies (AP).
Appendix C Implementation Consistent with Diffusion Policy
We adopt similar experimental settings with 1 or 2 observations, a prediction horizon of 16, and an executable action length of 8, following the standard setup used in Diffusion Policy [11].
It is important to note that our classical formulation introduces a minor discrepancy in the horizon definition compared to the implementation of Diffusion Policy [11]. Specifically, in our formulation, the horizon $H$ encompasses the past observed steps, meaning that the index of the current next predicted action is $O$, rather than 0. In contrast, as outlined in formulation Eq. 1, the horizon $H$ does not include past observations, with the first prediction step corresponding to the next time step. While the rationale behind this design remains unclear due to limitations in the author’s understanding, we retain the horizon definition introduced by Diffusion Policy [11] to ensure consistency in our experimental comparisons.
Figure 12: Visualization of Tasks in Single-Task Experiment.
Figure 13: Visualization of Tasks in Multi-Task Experiment.
Figure 14: Visualization of All Interaction Tasks in Kitchen Experiment.
Figure 15: Visualization of Tasks in Real-World Setup.